OpenStudy (anonymous):

find the general solution \[y'-8 y+18=0\qquad y(2)=1\]

OpenStudy (anonymous):

what's qquad

OpenStudy (anonymous):

\[y'-8 y+18=0\]

OpenStudy (anonymous):

forget it

OpenStudy (anonymous):

no idea what quad is

OpenStudy (anonymous):

nah, it just said it by your question in my browser

OpenStudy (anonymous):

really its an init val prob, but i can do that part.

OpenStudy (anonymous):

i didnt mean to copy the y(2) = 1

OpenStudy (anonymous):

what i got is \[ln(8y-18) = 8t + C\]

OpenStudy (anonymous):

but with the init value it gives an illegal ln value, ln(-10)

OpenStudy (anonymous):

\[y'-8y = -18 \Rightarrow e^{-8t}y'-8e^{-8t}y = -18e^{-8t}\] \[\Rightarrow (ye^{-8t})' = -18e^{-8t} \Rightarrow ye^{-8t} = \frac{-18}{-8}e^{-8t}+C\] \[y = \frac{9}{4}+Ce^{8t}\]

OpenStudy (anonymous):

The answer I got is y=9/4 - (5e^(8x))/(4e^(16))

OpenStudy (anonymous):

Now you use the initial value to find out C

OpenStudy (anonymous):

If there are any steps you need explained in more detail please ask :)

OpenStudy (anonymous):

Bingo. Thanks Dave

OpenStudy (anonymous):

No problem, you've helped me in the past ;-)

OpenStudy (anonymous):

I had something close but not quite: \[y(x)=(-5*e^{8(t+2)}+9)/4\]

OpenStudy (anonymous):

yeah, instead of the (t+2) it should be minus 2 since there is division there not multiplication